OpenStudy (anonymous):

If f(x)=3x+1 and f-1 (inverse)=x-1 over 2 Then f(2)= 1/3 1 7 please help with functions and relations:inverses!!

zepdrix (zepdrix):

\[\Large f(\color{royalblue}{x})=3(\color{royalblue}{x})+1\] \[\Large f(\color{royalblue}{2})=3(\color{royalblue}{2})+1\]

zepdrix (zepdrix):

Function notation can take a little while to get comfortable with :) Do the colors help at all?

zepdrix (zepdrix):

For f(2) we're letting x=2 in all the places we would normally have an x.

OpenStudy (anonymous):

@zepdrix yes it really does...what made me confuse is I thought I was suppose to plug it into the x-1 over two

zepdrix (zepdrix):

I'm not sure why they gave us that piece of information (about the inverse function). Are there other questions to follow that use these two functions maybe? Others asking about \(\large f^{-1}(2)\) or something maybe?

OpenStudy (anonymous):

yes I don't either, it really threw me off....so it was the same question how would I find f-1(7) same idea

zepdrix (zepdrix):

\[\Large f^{-1}(\color{teal}{x})\quad=\quad \frac{(\color{teal}{x})-1}{2}\] \[\Large f^{-1}(\color{teal}{-7})\quad=\quad \frac{(\color{teal}{-7})-1}{2}\]

zepdrix (zepdrix):

If this is supposed to be the inverse of the first f they provided, the denominator should be a 3, not a 2. Hmmm.

OpenStudy (anonymous):

Yes it there is a 2, I got it wrong because I had -2

OpenStudy (anonymous):

So (-7)-1 over two...I'm not understanding how we are suppose to get two...also another option I have is 22, 2, -2

zepdrix (zepdrix):

For the inverse question?\[\Large \frac{-7-1}{2}\quad=\quad \frac{-8}{2}\quad=\quad -4\]That's not one of the options?

zepdrix (zepdrix):

Oh lolol :) You asked for f^{1}(7)... not -7, sorry i misread that. :p

zepdrix (zepdrix):

\[\Large f^{-1}(\color{teal}{x})\quad=\quad \frac{(\color{teal}{x})-1}{2}\] \[\Large f^{-1}(\color{teal}{7})\quad=\quad \frac{(\color{teal}{7})-1}{2}\]

zepdrix (zepdrix):

So it looks like we're getting uhhh, 3 yes?\[\Large \frac{7-1}{2}\quad=\quad\frac{6}{2}\quad=\quad 3\]

OpenStudy (anonymous):

No I'm sorry there is not a three :/

zepdrix (zepdrix):

Can you take a picture of the problem maybe? D:

OpenStudy (anonymous):

If f(x)=3x+1 and f-1= x-1 over 3 then f -1(7) = 22, 2,-2 ???

zepdrix (zepdrix):

`Yes it there is a 2,`.... I thought u said the denominator was a 2...

OpenStudy (anonymous):

Oh...I'm sorry about that, it was my fault

zepdrix (zepdrix):

\[\Large f^{-1}(\color{teal}{x})\quad=\quad \frac{(\color{teal}{x})-1}{3}\] \[\Large f^{-1}(\color{teal}{7})\quad=\quad \frac{(\color{teal}{7})-1}{3}\]Ok if we have a 3, it should give us one of the answers on the list :D

OpenStudy (anonymous):

But yes, that makes a lot of sense the way you set it up...I really got that:)

zepdrix (zepdrix):

cool c:

OpenStudy (anonymous):

Yep thanks a lot, think I'm ready to take my quiz now:)

OpenStudy (anonymous):

Let R={0, 1, 2, 3} be the range of h(x) = x - 7. The domain of h inverse is {-7, -6, -5, -4} {7, 8, 9, 10} {0, 1, 2, 3} can you please explain this last one, I got the last answer and it was right but I just took a wild guess on it:/

zepdrix (zepdrix):

When we take the inverse of a function, the `domain` and `range` are exchanged. So if the `domain` of \(\large h(x) \)is R, then the `range` of \(\large h^{-1}(x)\) will also be R.

zepdrix (zepdrix):

Here's a way to think of it maybe, ummm.... If f(x) has `domain` D and `range` R, then f inverse will have `domain` R and `range` D.

OpenStudy (anonymous):

@zepdrix I understand that, but like how come the answer stay the same...Oh so the range would be (7, 8,9,10) then when you interchange them...it would be like 7-7=0, 8-7=1,9-7=2, and 10-7=3 so that the answer would just stay the same??? Did I understand that right?

zepdrix (zepdrix):

So for h(x): R={0, 1, 2, 3} If we plug these in for h(x) we get a domain of: D={-7, -6, -5, -4} ----------------------------------------- So for h(x): Domain={-7, -6, -5, -4} Range={0, 1, 2, 3} ----------------------------------------- For h inverse, we switch them: Domain={0, 1, 2, 3} Range={-7, -6, -5, -4}

zepdrix (zepdrix):

h(x) Range={0, 1, 2, 3} h inverse Domain={0, 1, 2, 3} right? :o

OpenStudy (anonymous):

@zepdrix Yes that's right:D and I get it now but I'll study more before I take my quiz...anyways thank you so much for the clarifications it helped a lot

zepdrix (zepdrix):

\c:/ np